# Surveying the Landscape of Image Captioning Evaluation: A Comprehensive Taxonomy and Novel Ensemble Method

### v1

@@ 5.3 Correlation with Human Ratings
Table 3 presents the results. We omit results for five metrics [^8] CLIPImageScore, Exact NO, Exact VO, Fuzzy NO, Fuzzy VO. that were proposed ad hoc in an experiment (rather than in a dedicated paper) and showed weak performance. We will now discuss the findings derived from this data.
@@
Our proposed EnsembEval achieves state-of-the-art results across all datasets, indicating that integrating various aspects of caption quality yields a more “human-like” score.
@@
The five dominant metrics, as well as their ensemble, exhibit weaker correlations with human ratings compared to other metrics. Moreover, BLEU4, the most common BLEU variation, shows the weakest correlation with human ratings.
@@ 5.4 Accuracy on Pairwise Comparison
The results are presented in Table 4. EnsembEval attained the highest score on the Reformulations dataset, as well as when comparing human-generated to machine-generated captions in Pascal 50 (HM). In all other cases MPNetScore and BLIP2Score performed best. As noted earlier, these metrics were absent from previous studies since they weren’t introduced in dedicated papers.
@@ 6.1 Image Captioning Surveys
A few studies Stefanini et al. (2022); Xu et al. (2023) provide a more detailed taxonomy of automatic metrics. Our work differs from them in two ways: First, we conduct a systematic review to gather all metric usage, identifying strong metrics not covered previously (as discussed in Section 5.3). Second, while these studies group metrics into loosely defined categories (e.g., “standard” metrics), we clearly define our categorization criteria based on the properties each metric aims to measure.
@@
Prior to our study, Staniūtė and Šešok (2019) reported metric usage from selected papers; however, they manually selected popular papers rather than identifying them systematically. Similar to our work, two studies Sharma (2021); Al-Shamayleh et al. (2024) systematically identify image captioning papers and report metric usage. However, they cover much fewer papers (79 and 41, respectively) and metrics (3 and 8, respectively), and do not report trends or provide in-depth analysis.
@@ 6.2 Image Captioning Metrics Analysis
Recently, as reference-free metrics like CLIPScore gain prominence, there has been growing criticism directed at this line of research. Deutsch et al. (2022) argue that reference-free metrics can be exploited at test time to find outputs that maximize their scores. Ahmadi and Agrawal (2024) find these metrics sensitive to visual grounding errors but not to caption implausibility. Ma et al. (2024b) show that optimizing for reference-free metrics leads to unreadable captions.
@@ 7 Conclusion
We conducted a comprehensive survey of image captioning evaluation methods, finding that the vast majority of papers use only 5 metrics which are weakly correlated with human judgment. Instead, we propose EnsembEval, which leverages a diverse set of metrics in a data-driven manner to achieve state-of-the-art results.
Acknowledgements
@@ A.1 Datasets
We use the official version of the Flickr8k dataset.999www.kaggle.com/datasets/dibyansudiptiman/flickr-8k We follow previous work Anderson et al. (2016); Hessel et al. (2021) when handling references that are also included in the candidate set: in Flickr8k-Expert we remove these sentences from the candidate set (158 examples), and in Flickr8k-CF we remove these sentences from the reference set.
@@
We use the official version of the Composite dataset.101010imagesdg.wordpress.com/image-to-scene-description-graph While previous work mentioned that there are three candidates per image and a single score per candidate, we find that the dataset contains four candidates per image for MSCOCO and Flickr30k, and two scores per candidate (correctness and throughness). We find that results from prior work was most closely replicated when we discarded the fourth candidate when one exists and used the correctness score as the total score. Also, following Hessel et al. (2021) we remove sentences that appear both as references and as candidates from the reference set.
@@
We use the official version of the THumB dataset.111111github.com/jungokasai/THumB Following the practice in other datasets, we remove sentences that appear both as references and as candidates from the reference set.
@@
We use the HuggingFace version of the Polaris dataset.121212huggingface.co/datasets/yuwd/Polaris
@@
The link to download the official version of the dataset in the CIDEr blog post131313vrama91.github.io/cider/ seems to be broken; we thank (REMOVED FOR ANONIMYZATION) for providing us with the original files.
@@
For the five most dominant metrics we use the pycocoevalcap implementation.141414pypi.org/project/pycocoevalcap
@@
We use an implementation151515github.com/DavidMChan/caption-by-committee/blob/main/cbc/metrics/content_score.py provided to us by the author.
@@
Since no implementation is provided by the authors, we implement the metric ourselves. We use the SD-XL 1.0-base model161616huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to generate the image based on the candidate.
@@
Since no implementation is provided by the authors, we implement the metric ourselves. We use the all-mpnet-base-v2 model.202020huggingface.co/sentence-transformers/all-mpnet-base-v2
@@ A.3 Feature Selection and Linear Regression
We use the scikit-learn212121scikit-learn.org/stable implementations of feature selection and linear regression. We use forward-selection in feature selection.
@@ Appendix B Metric Usage Records
Machine Learning Ma et al. (2024c): LexSim, RefPhSim, Black et al. (2024b): LexSim, RefSenSim, Liu et al. (2024): LexSim, RefPhSim, Ma et al. (2024a): LexSim, Qi et al. (2024): LexSim, Qiu et al. (2024): LexSim, RefPhSim, Wang et al. (2024): LexSim, RefPhSim, Fu et al. (2024): LexSim, RefPhSim

### v2

@@ 6.4 Correlation with Human Ratings
Table 4 presents the results. We omit results for five metrics [^12] CLIPImageScore, Exact NO, Exact VO, Fuzzy NO, Fuzzy VO. that were proposed ad hoc in an experiment (rather than in a dedicated paper) and showed weak performance.
@@
EnsembEval demonstrates the highest correlation with human ratings across all datasets, indicating that integrating various aspects of caption quality yields a more “human-like” score.
@@
The five dominant metrics, as well as their ensemble, correlate less with human ratings compared to all other examined metrics.
@@ 6.5 Accuracy on Pairwise Comparison
Results are presented in Table 5. EnsembEval attained the highest score on the Reformulations dataset, as well as when comparing human- and machine-generated captions in Pascal 50 (HM). In all other cases MPNetScore and BLIP2Score performed best. As noted earlier, these metrics were absent from previous studies since they were not introduced in dedicated papers.
6.6 Why Does the Ensemble Improve Correlation?
To understand why EnsembEval improves correlation over individual metrics, we manually examine cases where scores of individual metrics deviate from human ratings while the ensemble score aligns more closely.
Figure 6: An image, corresponding candidate caption, reference example and evaluation scores. Since no reference captions mention the blanket, Polos (a reference-based metric) assigns a low score. However, as the caption accurately describes the image, BLIP2Score (a reference-free metric) gives a high score. The ensemble score falls in between, aligning best with human ratings. Image and caption taken from the Polaris dataset.
These cases often arise from disagreements on which entities to describe. If reference captions omit an entity mentioned in the candidate caption, reference-based metrics assign a low score, while reference-free metrics give a high one. Humans typically rate such captions moderately, as they are accurate but may not align with their descriptive choices. The ensemble, combining both metric types, aligns more closely with human ratings. This issue is especially relevant beyond English, as research shows speakers of different languages tend to describe different entities Berger and Ponti (2024). See Figure 6 for an example.
@@ 7.1 Image Captioning Surveys
A few studies Stefanini et al. (2022); Xu et al. (2023) provide a more detailed taxonomy of automatic metrics. Our work differs from them in two ways: First, our systematic review identifies strong metrics not covered previously. Second, while these studies group metrics into loosely defined categories (e.g., “standard” metrics), we clearly define our categorization criteria based on the properties each metric aims to measure.
@@
Prior to our study, Staniūtė and Šešok (2019) reported metric usage from selected papers; however, they manually selected popular papers rather than identifying them systematically. Similar to our work, two studies Sharma (2021); Al-Shamayleh et al. (2024) systematically identify image captioning papers and report metric usage, but cover much fewer papers (79 and 41, respectively) and metrics (3 and 8, respectively), and do not report trends or provide in-depth analysis.
@@ 7.2 Image Captioning Metrics Analysis
Recently, as reference-free metrics like CLIPScore gain prominence, there has been growing criticism directed at this line of research. Deutsch et al. (2022) argue that reference-free metrics can be exploited at test time to find outputs that maximize their scores. Ahmadi and Agrawal (2024) find these metrics sensitive to visual grounding errors but not to caption implausibility.
@@ 8 Conclusion
We conducted a comprehensive survey of image captioning evaluation methods, yielding two key outcomes: a taxonomy of captioning metrics, including overlooked high-performing ones (e.g., BLIP2Score) and an analysis revealing that most papers rely on five metrics with weak human correlation. We further showed that an ensemble of metrics improves correlation with human ratings, highlighting the potential for further advancements in this approach.
In addition to the theoretical contribution, we anticipate two practical uses for EnsembEval. First, it is readily available131313github.com/uriberger/caption_evaluation as a reliable metric with strong correlation to human ratings. Second, users can extend the ensemble by incorporating additional metrics or adjusting the weights.
Acknowledgments
@@ A.1 Datasets
We use the official version of the Flickr8k dataset.141414www.kaggle.com/datasets/dibyansudiptiman/flickr-8k We follow previous work Anderson et al. (2016); Hessel et al. (2021) when handling references that are also included in the candidate set: in Flickr8k-Expert we remove these sentences from the candidate set (158 examples), and in Flickr8k-CF we remove these sentences from the reference set.
In Flickr8k-expert, annotators are instructed to rank the candidate on a scale of 1 to 4, with the following specifications:
4 = Sentence describes the image: Sentence contains no errors, everything described in the sentence appears in the image.
3 = Sentence almost describes the image: Sentence contains only a single moderate error or a small number of minor errors, major details described in the sentence appear in the image.
2 = Sentence barely describes the image: Sentence contains a major error or many moderate or minor errors, only some minor details described in the sentence appear in the image.
1 = Sentence does not describe the image at all: No details described in the sentence appear in the image, sentence is unrelated to image.
In Flickr8k-CF, annotators are shown images with ten sentences below each images, and are asked to select all sentences that could describe the image, with the following specifications:
Yes- The sentence is a good enough description of the image (minor details may be wrong, and not everything that appears in the image has to be described).
No- The sentence does not describe the image because it contains major errors or is completely unrelated to the image.
@@
We use the official version of the Composite dataset.151515imagesdg.wordpress.com/image-to-scene-description-graph While previous work mentioned that there are three candidates per image and a single score per candidate, we find that the dataset contains four candidates per image for MSCOCO and Flickr30k, and two scores per candidate (correctness and throughness). We find that results from prior work was most closely replicated when we discarded the fourth candidate when one exists and used the correctness score as the total score. Also, following Hessel et al. (2021) we remove sentences that appear both as references and as candidates from the reference set.
Instructions for the correctness rating (which we use in this study) are as follows:
Descriptions that correctly describe the image content with higher precision have better correctness ratings.
Score 1: The description has no relevance to the image.
Score 2: The description have only weak relevance to the image.
Score 3: The description have some relevance to the image.
Score 4: The description relates closely to the image.
Score 5: The description relates perfectly to the image.
@@
We use the official version of the THumB dataset.161616github.com/jungokasai/THumB Following the practice in other datasets, we remove sentences that appear both as references and as candidates from the reference set.
In this study, the authors performed the annotations themselves, relying on general descriptions for each aspect of the rating rather than providing specific instructions.
@@
We use the HuggingFace version of the Polaris dataset.171717huggingface.co/datasets/yuwd/Polaris
Annotators were instructed to rate the candidate on a scale of 1 to 5, with the following specifications:
Evaluate the extent to which the sentences capture the features of the image, considering fluency, relevance and descriptiveness.
Use a five-point scale for your evaluation, employing a point addition or deduction system.
For descriptiveness, rate a caption as ‘Excellent’ if it is both detailed and accurate.
For relevance, award points if the caption accurately depicts any present object, even if irrelevant words are included. Deduct points for each irrelevant word.
For fluency, assign lower scores if there are grammatical errors.
For example, consider the caption ‘A man wearing a helmet is skateboarding with a dog behind him.’ with the given image (Figure 7). This caption scores high in relevance as it accurately describes the main object in the image. However, the mention of ‘a dog,’ which is absent from the image, reduces its descriptiveness. Despite its high fluency, this warrants a ’Fair’ (3 on a scale of 1 to 5) rating overall.
Figure 7: Image provided as an example in the Polaris dataset instructions.
@@
The link to download the official version of the dataset in the CIDEr blog post [Link] seems to be broken; we thank (REMOVED FOR ANONIMYZATION) for providing us with the original files.
@@
For the five most dominant metrics we use the pycocoevalcap implementation. [Link]
@@
We use an implementation [Link] provided to us by the author.
@@
Since no implementation is provided by the authors, we implement the metric ourselves. We use the SD-XL 1.0-base model [Link]] to generate the image based on the candidate.
@@
Since no implementation is provided by the authors, we implement the metric ourselves. We use the all-mpnet-base-v2 model. [Link]
@@ A.3 Feature Selection and Linear Regression
We use the scikit-learn [Link] implementations of feature selection and linear regression. We use forward-selection in feature selection.
@@ Appendix B Metric Usage Records
Machine Learning Ma et al. (2024b): LexSim, RefPhSim, Black et al. (2024b): LexSim, RefSenSim, Liu et al. (2024): LexSim, RefPhSim, Ma et al. (2024a): LexSim, Qi et al. (2024): LexSim, Qiu et al. (2024): LexSim, RefPhSim, Wang et al. (2024): LexSim, RefPhSim, Fu et al. (2024): LexSim, RefPhSim